OpenStudy (anonymous):

Solve the equation. x^2+18x+81=25

OpenStudy (anonymous):

Where do you have problems there?

OpenStudy (anonymous):

that is the problem

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

So is the first quadratic equation that you meet?

OpenStudy (anonymous):

@Grazes i need help @Marrioncito i dont get what you saying

OpenStudy (anonymous):

Is this the first time you have to solve a quadratic equation?

OpenStudy (anonymous):

@Marrioncito yes i have solved them before.. but im asking.. how do you solve it?

OpenStudy (anonymous):

Yo have to use the "quadratic formula". In order to apply it, you have to write the polinomia. Like any other quadratic ecuation.

OpenStudy (anonymous):

\[x^2+18x+81 - 25 = 0 = x^2+18x+56\] Then, a = 1, b = 18, and c = 56...and apply the quadratic formula.

OpenStudy (anonymous):

Have you ever used the quadratic formula?

OpenStudy (anonymous):

(x+9)^2=25

OpenStudy (anonymous):

sqrt(25=5 5=x+9 x=-4

OpenStudy (anonymous):

@Marrioncito no i have not.. @Grazes lol.. please leave him alone and I do not know which one of you all suggestions I should take

OpenStudy (anonymous):

Have you learned Squaring the Root yet? Or just factoring?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

So you just square the root and factor the one side because it is already a perfect square. You get x+9 to the power of two. Square root the 25 on the other side as if it were a regular two sided equation. You get x+9=5. 5-9=x. x=-4

OpenStudy (anonymous):

Better?

OpenStudy (anonymous):

@ Grazes you gotta b wrong because i have multiple choice and it does not have a 5 in it.. but it have -4.. you not calculating something right..???

OpenStudy (anonymous):

Is it wrong. Do you want the method or just you want the results? x = -4 , x = -14

OpenStudy (anonymous):

The answer is -4

OpenStudy (anonymous):

where did you get 5 o-o

OpenStudy (anonymous):

i want the method as well not just the answers.. @ Grazes its two answers though

OpenStudy (anonymous):

@Marrioncito i want the method to, pleasee

OpenStudy (anonymous):

See quadratic formula; http://en.wikipedia.org/wiki/Quadratic_equation#Quadratic_formula

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Then read the message that I wrote ten minutes ago.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

And it as a whole ll make sense.

OpenStudy (anonymous):

If something isn´t clear, just ask! =)

OpenStudy (anonymous):

lol. i got it.. thanks Marrioncito!!

OpenStudy (anonymous):

You´re welcome

OpenStudy (anonymous):

btw, for mine, one of the steps should've been -5=x+9 which would make x=-14. This is because the square root of 25 is positive OR negative 5